If Leah is 6 years older than her sister, Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?
Found 2 solutions by ankor@dixie-net.com, fama:
Answer by ankor@dixie-net.com(22740)

Write a simple equation for each sentence
:
If Leah is 6 years older than her sister, Sue,
L = S + 6
:
John is 5 years older than Leah,
J = L + 5
:
the total of their ages is 41.
L + S + J = 41
:
Then how old is Sue?
Take the 1st equation, rearrange
S + 6 = L
S = (L - 6)
:
Substitute (L-6) for S and (L+5) for J in the "total age" equation, find L:
L + (L-6) + (L+5) = 45
3L - 6 + 5 = 41
3L - 1 = 41
3L = 41 + 1
L = 42/3
L = 14 yrs old is Leah
:
I'll let you find ages of S & J; check your solution in the total age equation

Answer by fama(1)